Find the point of intersection and the angle between 2x-y+3=0 and 3x+y-1=0. Use decimal values for fractional answers.
Answer by Alan3354(69443)

2x-y+3=0 and 3x+y-1=0
-----------
2x-y+3=0
3x+y-1=0
----------- Add them